Co-60 is used extensively in radiotherapy as a source of y-rays. Its half-life is 5.27 years. a) How many years will it take to decrease to 35% of its original activity? b) If a cobalt source is ten years old, what percent of its original activity remains?


1
Expert's answer
2021-03-12T06:29:33-0500

a) "\lambda" = 1/t1/2 = 1/5.27 y-1

t =( 1/"\lambda" ) ln "\frac{N_o}{N_t}"

decrease to 35% of No

t = 5.27 × ln "\frac{N_o}{0.35N_o}"

t = 5.27× ln( 2.857)

t = 5.27×1.04972

t = 6.004 years

b) Nt = No × e(--1/5.27)×10

Nt = No × e-1.89

Nt = 0.15No

15%